8-pack control system working group meeting
4 December 2002
In attendence: Chestnut, Clark, Kim, Nelson, S Smith, Young
Notes by: J Nelson
Generic agenda was status of the following:
LLRF SW module support
Crates and requirement changes
Klystron/structure processing loop
Janice reported that the PLC checklist is still in the creation process. Hopes to have it by the end of the week. Judy comes back on Monday. In the 8-pack meeting yesterday Dick made some statements that implied that his code has changed somewhat. Ron/Spence will start the process of asking for the latest version of his variable tables and touch panels.
Ron reported that the ADC database is coming soon – in the next few days.
Andrew said the ADC, DAC, SIS and DIO modules are ready. IQA and VSTEP ready next week.
Andrew reported that some of the cabling for the ADC (big fire hose) is not quite ready, but the rest is ready.
Andrew will provide test cables to SW for testing the ADCs and DACs as well as the DIO fanout board.
Kukhee said the trigger module for the ADCs is working and is on his desk.
Ron and Kukhee reported the SIS software is ready. The diagnostic displays work, but not the normal displays. After the meeting, we all looked at the buttons to access the waveform PVs and they were not connected. Kukhee and Ron need to resolve the outstanding SIS database issues.
The waveform archiving works. Accessing the desired waveform from the archiver database Ron said would be challenging. Janice explained their initial anticipated use (acoustic sensor breakdown analysis) and was reassured by Ron and Spence that pulling things from the archiver to matlab is ready to go.
Janice will research what 5-th order polynomial calibrations are needed for 8-pack as she’s forgotten.
Ron will pull together a list of the remaining database work.
The klystron/structure processing loops are next on the list after the ADC, DAC, SIS, DIO, etc software is working. Steve warned that the pulse width control will be different from NLCTA’s.
Ron reported that the crate profiles on the web are out of date. He and Kukhee will verify with hardware in ESB then update the web and (maybe?) print out what it’s supposed to look like and tape it to the racks.
After the meeting, we all looked at the existing LLRF epics panels. Obvious problems not mentioned above:
The waveform acquire button is still white. Needs a database and maybe more? This is for Ron/Kukhee/Judy.
For Janice, check to make sure the SCP config templates are up to date.
For Janice, fix the SLED panel falling off the bottom of the screen.
For Ron, et al, the DIO database doesn’t seem to exist yet either.
Other outstanding issues that Janice thought up after the meeting:
For Ron et al, does the PV database exist for 8-pack Peak Power Meters?